If tan^2a=1+2tan^2b,then prove that cos 2b=1+2cos2a
1
Expert's answer
2019-05-02T10:11:20-0400

The identity: tan2a = "\frac{1}{cos^{2}a}" - 1; tan2b = "\frac{1}{cos^{2}b}" - 1;

"\implies" "\frac{1}{cos^{2}a}" - 1 = 1 + 2("\frac{1}{cos^{2}b}" - 1);

"\implies" "\frac{1}{cos^{2}a}" = "\frac{2}{cos^{2}b}" ; cos2b = 2cos2a;

The identity: cos2b = "\frac{1+cos2b}{2}"; cos2a = "\frac{1+cos2a}{2}" ;

"\implies" "\frac{1+cos2b}{2}" = 1 + cos2a;

"\implies" cos2b = 1 + 2cos2a.

Answer: from the relation tan2a = 1 + 2tan2b, follows: cos2b = 1 + 2cos2a.